OpenStudy (anonymous):

Factor the trinomial: 3x2 + 11x + 6 A. (3x + 2)(x + 3) B. (3x + 3)(x + 2) C. (3x + 1)(x + 6) D. (3x + 6)(x + 1)

OpenStudy (hoblos):

3x2 + 11x + 6 = 3x2+ 2x + 9x +6 = x(3x+2) + 3(3x+2) = (3x+2)(x+3)

OpenStudy (anonymous):

A